# references.bib

### v5

@@ Appendix A Genetic programming
1:initialise population $P$ with size $N$ (Section 3.4.1)
@@
3:     evaluate each individual in $P$ on $F$ (Section 3.4.2)
@@
9:         children $c$ = reproduce($p$) (Section 3.4.3)
@@
Hyperparameter Exp 1 Exp 2 Exp 3 Exp 1 Exp 2 Exp 1
Population Size 500 500 1000 500 500 1000
Number of Generations 30 50 100 50 50 100
Latent state size 2 2 2 2 2 2
@@ Appendix B Stochastic Harmonic Oscillator
GP and the baselines are tested on the SHO in three different experiments. In the first experiment the policies interact with the standard setting of the SHO. In this experiment $\mathbf{A}$ is specified with $\omega=1$ and $\zeta=0$ and the state observations are noisy. In the second experiment the state of the SHO is partially observed, as the policies only observe the position while the velocity is unknown. $\mathbf{A}$ remains the same in this experiment. In the third experiment, the policies are evaluated on generalisability. In each trial $\omega$ and $\zeta$ are different, specifically $\omega\sim U(0,2)$ and $\zeta\sim U(0,1.5)$. In this experiment both the position and velocity are observed again. In the experiments with observation noise, the parameters of the readout function are set to $\mathbf{D}=\mathbf{I}$ and $\mathbf{\Sigma}=0.3\mathbf{I}$ (see Equation 2). In experiments with partial state observability, $\mathbf{D}$ is adapted such that only the relevant observations are returned and the dimension of $\mathbf{\Sigma}$ is decreased to match the number of observations.
@@ Appendix C Acrobot
In every experiment, $\mathbf{R}\in\mathbb{R}^{C\times C}$ is set to $0.01\mathbf{I}$, $x^{\ast}$ to 1.5 and the initial states are sampled from $U(-0.1,0.1)$. Observations of $\theta_{i}$ are wrapped between [-$\pi$, $\pi$] and a trial terminates when $\dot{\theta_{1}}>8\pi$ or $\dot{\theta_{2}}>18\pi$. Policies are discovered to solve the acrobot in three different experiments. In the first experiment the state of the acrobot is observable with noise and control force is only applied to the second link. In the second experiment the acrobot state is made partially observable, as the policies no longer observe the angular velocities. In the third experiment, the policies can apply independent forces to both links, and the angular velocities are observed again. The observations of the acrobot include angles, therefore the sine and cosine functions are added to the function set of the GP algorithm, besides the arithmetic operators. As for the SHO, the parameters of the readout function are set to $\mathbf{D}=\mathbf{I}$ and $\mathbf{\Sigma}=0.3\mathbf{I}$ to add observation noise. In experiment with partial state observability, $\mathbf{D}$ is changed to exclude the unobserved variables and $\mathbf{\Sigma}$ is set to the correct dimension.
@@
Appendix D Continuous stirred tank reactor
$d\mathbf{x}=f(\mathbf{x},\mathbf{u})dt+\mathbf{V}d\mathbf{w}$ (20)
with
where
$\mathbf{Q}=\begin{bmatrix}0&0&0\\ 0&0.01&0\\ 0&0&0\end{bmatrix},\quad r=0.0001,\quad\mathbf{x}^{\ast}=\begin{bmatrix}0\\ T_{r}^{\ast}\\ 0\end{bmatrix}.$ (26)
Only a single experiment is performed regarding the CSTR. This experiment tests policies for their generalisability to an industrial environment. The parameters of the CSTR are different in each trial, to simulate as if the policy is operating in multiple scenarios. In total, eight parameters vary between trials (Table III). Additionally, the concentration in the reactor is not observed, and the observations of the temperature in the reactor and cooling jacket are noisy. The function set of the GP algorithm is extended to include the exponential and logarithmic functions to aid the policies in dealing with the values of the temperatures, which typically range between 200 and 600. To generate the observations, $\mathbf{D}$ is defined such that the concentration is excluded from the observations and $\mathbf{\Sigma}=7.5\mathbf{I}$.

### v6

@@ Appendix A Genetic programming
1:initialise population $P$ with size $N$ (Section III-D1)
@@
3:     evaluate each individual in $P$ on $F$ (Section III-D2)
@@
9:         children $c$ = reproduce($p$) (Section III-D3)
@@
Hyperparameter Exp 1 Exp 2 Exp 3 Exp 1 Exp 2 Exp 3 Exp 1
Population Size 500 500 1000 500 500 500 1000
Number of Generations 30 50 100 50 50 50 100
Latent state size 2 2 2 2 2 3 2
@@ Appendix B Stochastic Harmonic Oscillator
GP and the baselines are tested on the SHO in three different experiments. In the first experiment the policies interact with the standard setting of the SHO. In this experiment $\mathbf{A}$ is specified with $\omega=1$ and $\zeta=0$ and the state observations are noisy. In the second experiment the state of the SHO is partially observed, as the policies only observe the position while the velocity is unknown. $\mathbf{A}$ remains the same in this experiment. In the third experiment, the policies are evaluated on generalisability. In each trial $\omega$ and $\zeta$ are different, specifically $\omega\sim U(0,2)$ and $\zeta\sim U(0,1.5)$. In this experiment both the position and velocity are observed again. In the experiments with observation noise, the parameters of the readout function are set to $\mathbf{D}=\mathbf{I}$ and $\mathbf{\Sigma}=\sigma\mathbf{I}$ with $\sigma=0.3$ (see Equation 2). In experiments with partial state observability, $\mathbf{D}$ is adapted such that only the relevant observations are returned and the dimension of $\mathbf{\Sigma}$ is decreased to match the number of observations.
@@ Appendix C Acrobot
In every experiment, $\mathbf{R}\in\mathbb{R}^{C\times C}$ is set to $0.01\mathbf{I}$, $x^{\ast}$ to 1.5 and the initial states are sampled from $U(-0.1,0.1)$. Observations of $\theta_{i}$ are wrapped between [-$\pi$, $\pi$] and a trial terminates when $\dot{\theta_{1}}>8\pi$ or $\dot{\theta_{2}}>18\pi$. Policies are discovered to solve the acrobot in three different experiments. In the first experiment the state of the acrobot is observable with noise and control force is only applied to the second link. In the second experiment the acrobot state is made partially observable, as the policies no longer observe the angular velocities. In the third experiment, the policies can apply independent forces to both links, and the angular velocities are observed again. The observations of the acrobot include angles, therefore the sine and cosine functions are added to the function set of the GP algorithm, besides the arithmetic operators. As for the SHO, the parameters of the readout function are set to $\mathbf{D}=\mathbf{I}$ and $\mathbf{\Sigma}=\sigma\mathbf{I}$ with $\sigma=0.3$ to add observation noise. In experiment with partial state observability, $\mathbf{D}$ is changed to exclude the unobserved variables and $\mathbf{\Sigma}$ is set to the correct dimension.
@@
Appendix D Continuous stirred tank reactor
[…]
with
where
$\mathbf{Q}=\begin{pmatrix}0&0&0\\ 0&0.01&0\\ 0&0&0\end{pmatrix},\quad r=0.0001,\quad\mathbf{x}^{\ast}=\begin{pmatrix}0\\ T_{r}^{\ast}\\ 0\end{pmatrix}.$ (26)
Only a single experiment is performed regarding the CSTR. This experiment tests policies for their generalisability to an industrial environment. The parameters of the CSTR are different in each trial, to simulate as if the policy is operating in multiple scenarios. In total, eight parameters vary between trials (Table III). Additionally, the concentration in the reactor is not observed, and the observations of the temperature in the reactor and cooling jacket are noisy. The function set of the GP algorithm is extended to include the exponential and logarithmic functions to aid the policies in dealing with the values of the temperatures, which typically range between 200 and 600. To generate the observations, $\mathbf{D}$ is defined such that the concentration is excluded from the observations and $\mathbf{\Sigma}=\sigma\mathbf{I}$ with $\sigma=7.5$.
Appendix E Comparison with time-lagged policies
Table 6: The time-lagged static and dynamic symbolic policies discovered with genetic programming for controlling the partially observable harmonic oscillator with various levels of noise ($\sigma$). $u_{j}$ represents the function that outputs the control signal. In the time-lagged policies, the second index in the observations $y_{i},t$ indicates the delay. In the dynamic policies, the functions of $a_{i}$ represent the state equations of the latent variables. When a latent variable is not referred to in any of the other expressions, it is labelled as inactive.
$\sigma$ Static policy Time-lagged policy Dynamic policy
0.0 $\begin{array}[]{lll}u_{1}&=&0.25y_{1}+0.48x^{*}\\ &&-0.10\end{array}$ $\begin{array}[]{lll}u_{1}&=&35.44(y_{1,-1}-y_{1,-2})\\ &&+x^{*}\end{array}$ $\begin{array}[]{lll}u_{1}&=&a_{1}+x^{*}\\ \dot{a}_{1}&=&-2a_{1}+a2-x^{*}-0.18\\ \dot{a}_{2}&=&y_{1}-2u_{1}+x^{*}\end{array}$
0.25 $\begin{array}[]{lll}u_{1}&=&0.25y_{1}+0.49x^{*}\end{array}$ $\begin{array}[]{lll}u_{1}&=&3y_{1,-1}-y_{1,-2}-2y_{1,-3}\\ &&+0.97x^{*}\end{array}$ $\begin{array}[]{lll}u_{1}&=&1.40a_{1}+0.90a_{2}+x^{*}\\ \dot{a}_{1}&=&a_{2}+y_{1}+x^{*}\\ \dot{a}_{2}&=&-1.97u_{1}+y_{1}+x^{*}\end{array}$
0.5 $\begin{array}[]{lll}u_{1}&=&0.26y_{1}+0.47x^{*}\\ &&-0.05\end{array}$ $\begin{array}[]{lll}u_{1}&=&2y_{1,-1}-2y_{1,-3}+x^{*}\end{array}$ $\begin{array}[]{lll}u_{1}&=&a_{1}+x^{*}\\ \dot{a}_{1}&=&-0.72a_{1}+0.57y_{1}-0.57x^{*}\\ \dot{a}_{2}&=&0.14\,\text{(Inactive)}\end{array}$
0.75 $\begin{array}[]{lll}u_{1}&=&0.23y_{1}+0.52x^{*}\end{array}$ $\begin{array}[]{lll}u_{1}&=&1.18y_{1,-1}-0.91y_{1,-3}\\ &&+0.62x^{*}\end{array}$ $\begin{array}[]{lll}u_{1}&=&a_{2}+x^{*}\\ \dot{a}_{1}&=&-1.53(a_{1}+y_{1})\\ \dot{a}_{2}&=&-a_{1}-0.43a_{2}-u_{1}-0.01\end{array}$
1.0 $\begin{array}[]{lll}u_{1}&=&0.25y_{1}+0.47x^{*}\\ &&-0.05\end{array}$ $\begin{array}[]{lll}u_{1}&=&y_{1,-1}-0.77y_{1,-3}+0.66x^{*}\\ &&-0.06\end{array}$ $\begin{array}[]{lll}u_{1}&=&2.53a_{1}+3.53a_{2}\\ \dot{a}_{1}&=&-2a_{1}-y_{1}+x^{*}\\ \dot{a}_{2}&=&-a_{1}-u_{1}+x^{*}\end{array}$
Appendix F Variance Analysis
Figure 7: Comparison of fitness on training and testing data. (a) The fitness obtained by the best dynamic symbolic policy found during evolution on the stochastic harmonic oscillator (SHO) task with varying environment parameters is plotted. Additionally, the fitness of these policies obtained on unseen test data. The colour indicates the number of non-linear operations, counting multiplications between pairs of variables and the occurrences of divisions and powers. (b) The train and test fitness of the best dynamic symbolic policies from the continuous stirred tank reactor (CSTR) experiment. The non-linear operations can also include exponentials and logarithms here.